8. Explain briefly the working principle of a refrigerator and obtain anexpression for its coefficient of performance.

Grade:11

1 Answers

Arun
6 years ago
Dear Vaibhav
 
Refrigerator is a reverse heat engine.
Normally heat flows from a region of higher temperature to a region of lower temperature, and this heat can be converted to work. Thus the heat engine does work. This is the concept of heat engine.
Now if this be reversed like we have a machine that takes heat from the sink (QC) kept at temperature T2 and returns to the source (QH) kept at temperature T1 thereby making source cooler when work is done (W) on it. Thus in a refrigerator work is done by some external agency W to make source cooler.
Coefficient of performance (COP) of a refrigerator is defined as the ratio of the amount of heat drawn from the sink (QC ) to the work input W.
COP = QC/W
Work done can be calculated from 1st law of thermodynamics as
W = QH - QC
COP = QC/ (QH - QC)
= T2/(T1 – T2)  (From Carnot’s engine)

Regards
Arun (askIITians forum expert)